# Differentially private and decentralized randomized power method

### v1

@@ Abstract
The randomized power method has gained significant interest due to its simplicity and efficient handling of large-scale spectral analysis and recommendation tasks. As modern datasets contain sensitive private information, we need to give formal guarantees on the possible privacy leaks caused by this method. This paper focuses on enhancing privacy-preserving variants of the method. We propose a strategy to reduce the variance of the noise introduced to achieve Differential Privacy (DP). We also adapt the method to a decentralized framework with a low computational and communication overhead, while preserving the accuracy. We leverage Secure Aggregation (a form of Multi-Party Computation) to allow the algorithm to perform computations using data distributed among multiple users or devices, without revealing individual data. We show that it is possible to use a noise scale in the decentralized setting that is similar to the one in the centralized setting. We improve upon existing convergence bounds for both the centralized and decentralized versions. The proposed method is especially relevant for decentralized applications such as distributed recommender systems, where privacy concerns are paramount.
@@ 1 Introduction
The randomized power method has attracted significant attention due to its simplicity and its efficiency in solving large-scale linear algebra problems in spectral analysis or recommendation. It is in particular useful for computing the principal singular vectors of a matrix. It has a straightforward implementation which runs with low computational overhead (nearly linear complexity), making it particularly suitable for analyzing large modern datasets.
@@
To address and quantify privacy leakages, Differential Privacy (DP) has emerged as a powerful mathematical framework that provides formal guarantees and helps to mitigate potential privacy leaks of an algorithm. DP ensures that the output of an algorithm reveals little about any individual record in the input, even in the worst case. Several works have attempted to apply DP to the randomized power method. For instance, (Hardt & Price, 2014) and (Balcan et al., 2016) introduced differentially private versions of the method, and (Guo et al., 2021; Wang & Xu, 2020) explores federated approaches to private power iteration. Other close works explore centralized differentially private versions of PCA (Liu et al., 2022) and federated (Wang & Xu, 2020; Briguglio et al., 2023) and give optimal convergence bounds given certain assumptions.
@@ 1.1 Contributions
We propose an improved differentially private version of the randomized power method, with tighter convergence bounds, reducing the noise required to achieve privacy and therefore improving the accuracy of the method. In the process of deriving these bounds, we present a new, more straightforward proof of a privacy result for the differentially private randomized power method. [^1] This proof addresses some minor mistakes in a privacy proof from (Hardt & Roth, 2012, 2013) that have been reproduced in several follow-up works (Hardt & Price, 2014; Balcan et al., 2016). The modified result allows us to use a wider range of privacy parameters.
@@
Our contributions can be summarized as follows:
  1. 1.
    We introduce a novel privacy-preserving randomized power method that improves upon existing convergence bounds, with a reduced sensitivity to the number of singular vectors.
  2. 2.
    We extend the method to a decentralized setting, utilizing Secure Aggregation to ensure privacy in a distributed environment without the need for a trusted curator.
  3. 3.
    We provide new privacy proofs that correct errors in previous works, ensuring rigorous privacy guarantees.
  4. 4.
    We propose a new adjacency model to guarantee the differential privacy of our method, to open the door to new and more realistic use cases.
  5. 5.
    We present both runtime-dependent and runtime-independent bounds for the proposed methods, demonstrating their utility through theoretical analysis and empirical validation.
@@ 2.2.1 Approximate differential privacy:
A randomized algorithm $\mathcal{M}$ provides ($\epsilon,\delta$)-differential privacy if for all datasets $D_{1}$ and $D_{2}$ differing on at most one element (neighboring datasets), and for all measurable subsets $S\subseteq\text{Range}(\mathcal{M})$, the following holds:
@@ 2.2.2 $\ell_{2}$-sensitivity:
where $\|\cdot\|_{2}$ represents the $\ell_{2}$-norm. This sensitivity quantifies the stability of the function under small changes in the input data and allows us to calibrate the noise needed to ensure differential privacy. The notion of adjacency of the datasets is in general application-dependent. Its granularity reflects what one wants to protect by using Differential Privacy. In (Hardt & Price, 2014; Balcan et al., 2016), it is defined as a magnitude 1 one-element change in the matrix of interest. In the context of recommender system, this would mean that we want to protect individual user-item interactions.
@@ 2.2.3 Zero-Concentrated Differential Privacy (zCDP):
where $D_{\alpha}$ is the Rényi divergence of order $\alpha$ and $\rho$ is a positive parameter controlling the trade-off between privacy and accuracy (smaller values of $\rho$ imply stronger privacy guarantees). zCDP ensures that the privacy loss is tightly bounded in terms of the Rényi divergence, providing a more refined privacy analysis compared to pure or approximate differential privacy. Additionally, zCDP composes more easily than other differential privacy definitions and will be useful for our analysis.
@@ 2.3 Privacy preserving power iteration:
Let $k$ be a target rank and $\eta$ a small real controlling the approximation precision. The aim of the randomized power method is to find a matrix $\bm{Q}\in\mathbb{R}^{n\times k}$ such that $\|\bm{U}-\bm{Q}\bm{Q}^{\top}\bm{U}\|\leq\eta$. In other words, we want $\bm{Q}$ to approximate the $k$ principal eigenvectors of $\bm{A}$. The aim of the privacy-preserving randomized power method is to compute $\bm{Q}$ without leaking sensitive information embedded in $\bm{A}$. In (Hardt & Price, 2014; Balcan et al., 2016), a sensitive change is modeled as a single-element change of maximum Frobenius norm 1 in $\bm{A}$. According to (Hardt & Price, 2014; Balcan et al., 2016), given approximate-DP parameters ($\epsilon,\delta$) we can compute $\bm{Q}$ using Algorithm 1 with $\Delta_{l}=\Delta_{l}^{a}\triangleq\sqrt{p}\cdot\|\bm{X}_{l}\|_{\max}$.
@@
In recent works on the privacy-preserving power iteration (Hardt & Price, 2014; Balcan et al., 2016; Guo et al., 2021), the standard deviation of the noise related to Differential Privacy (DP) at iteration $l$ is:
$=\Delta_{l}^{a}\cdot\epsilon^{-1}\sqrt{4L\log(1/\delta)},$ (7)
as defined in Fig. 3 of (Hardt & Price, 2014), Alg. 2 of (Balcan et al., 2016) and Theorem 6 of (Guo et al., 2021). $\Delta_{l}^{a}$ is an $\ell_{2}$-sensitivity bound of the change in $\bm{A}\bm{X}^{\ell-1}$ given a change of magnitude at most 1 in $\bm{A}$ on only one element ($\Delta_{l}\leq\Delta_{l}^{a}$). Our first contribution comes from the fact that we can find a significantly tighter bound on $\Delta_{l}$.
@@ 2.4 Existing convergence bounds
To our knowledge, the strongest bound for the randomized power method as defined in Algorithm 1 with $\Delta_{l}=\Delta_{l}^{a}\triangleq\sqrt{p}\cdot\|\bm{X}_{l}\|_{\max}$ is given in (Balcan et al., 2016). In Theorem 2.4, we present a reduction of their Theorem 3.1. The theorem, and several others we present relies on the verification of the conditions in Assumption 2.3.
@@
Let $\bm{A}\in\mathbb{R}^{d\times d}$ be a symmetric data matrix. Fix target rank $k$, intermediate rank $q\geq k$, and iteration rank $p$, with $2q\leq p\leq d$. Suppose the number of iterations $L$ is set as $L=\Theta(\frac{\sigma_{k}}{\sigma_{k}-\sigma_{q+1}}\log(d))$. Let $\epsilon,\delta\in(0,1)$ be the differential privacy parameters. Let $\bm{U}_{q}\in\mathbb{R}^{d\times q}$ be the top-$q$ eigenvectors of $\bm{A}$ and let $\sigma_{1}\geq\cdots\geq\sigma_{d}\geq 0$ denote its eigenvalues. Then if Assumption 2.3 holds, Algorithm 1 with $\Delta_{l}=\Delta_{l}^{a}$ gives us the following runtime-dependent convergence bound:
@@ 3.1.1 Adjacency notion:
In (Hardt & Price, 2014; Balcan et al., 2016), a matrix $\bm{A}^{\prime}$ is denoted as adjacent to $\bm{A}$ ($\bm{A}\sim\bm{A}^{\prime}$) when $\bm{A}^{\prime}$ has a change of a single element of Frobenius norm at most $1$ compared to $\bm{A}$. In other words, we can write:
$\bm{A}^{\prime}=\bm{A}+c\cdot\bm{e}_{i}\cdot\bm{e}_{j}^{\top},$ (11)
with $\bm{e}_{i},\bm{e}_{j}\in\mathbb{R}^{d}$ being the canonical basis vector of dimension $d$. Without loss of generality, we have $0\leq i,j<n$ (the change of magnitude $c\leq 1$ can happen on any element $\bm{A}_{ij}$). In terms of differential privacy, this concept of adjacency allows us to protect element-wise updates to $\bm{A}$. We however note that as $\bm{A}$ and $\bm{A}^{\prime}$ are real and positive semi-definite, the change should also need to be on the diagonal of $\bm{A}$. We propose a new, more general, notion of adjacency to allow for other types of updates:
@@
with $\bm{C}$ a symmetric matrix representing the update, subject to $\sqrt{\sum_{i}\|C_{i:}\|_{1}^{2}}\leq 1$. The proposed adjacency notion is strictly more general than the previous adjacency notion (11). Indeed, $\bm{C}=c\cdot\bm{e}_{i}\cdot\bm{e}_{j}^{\top}$ is a special case. $\bm{C}$ can also for instance be non-zero on the diagonal or antidiagonal, or be another symmetric matrix.
@@ 3.1.2 Improved sensitivity bound:
The previous sensitivity bound $\Delta_{l}^{a}$ (used in (Balcan et al., 2016; Hardt & Price, 2014)) was found by using a straightforward extrapolation from the sensitivity of the change in $\bm{A}{\bm{X}^{\ell-1}}$ when ${\bm{X}^{\ell-1}}\in\mathbb{R}^{n\times 1}$ to when ${\bm{X}^{\ell-1}}\in\mathbb{R}^{n\times p}$. We can however find a tighter bound on $\Delta_{l}$ if we directly analyse the change in ${\bm{X}^{\ell-1}}\in\mathbb{R}^{n\times p}$. We give the result of our improved analysis using our proposed adjacency notion (12) in Theorem 3.1 and provide its proof in Appendix A.1.
@@
Let $\Delta_{l}=\|\bm{A}^{\prime}\bm{X}^{l}-\bm{A}\bm{X}^{l}\|_{F}$ where $\bm{A}^{\prime}$ is defined using (12). Then $\Delta_{l}$ can be upped bounded as follows:
$\begin{split}\Delta_{l}\leq\max_{i}\|\bm{X}_{i:}^{l}\|_{2}\triangleq\Delta_{l}^{b}\,.\end{split}$ (13)
3.1.3 Note on related privacy proofs:
Several Differential Privacy (DP) proofs for the private randomized power method (PPM) have been developed in prior works, for instance those by (Hardt & Price, 2014), (Balcan et al., 2016), (Hardt & Roth, 2013), (Hardt & Roth, 2012). These papers mostly rely on the same proofs to establish that the PPM satisfies ($\epsilon$, $\delta$)-DP. Specifically, (Balcan et al., 2016) references the privacy proof from (Hardt & Price, 2014), which builds upon the privacy proof from (Hardt & Roth, 2013). The proof therein is also closely related to the one of (Hardt & Roth, 2012).
However, both (Hardt & Roth, 2013) (Theorem 2.4) and (Hardt & Roth, 2012) (Theorem 2.4) contain errors in their proposed composition rules, where a comparison sign is mistakenly flipped. This error could potentially cause the privacy parameter $\epsilon$ in the proposed mechanism to be arbitrarily small, providing no privacy guarantee at all. The original composition rule is presented in Theorem III.3 of (Dwork et al., 2010). Moreover, Lemma 3.4 of (Hardt & Roth, 2013) misuses Theorem 2.4. Indeed, they claim that their algorithm satisfies $(\epsilon^{\prime},\delta)$-DP at each iteration. By their Theorem 2.4, then the algorithm overall should satisfy $(\epsilon^{\prime},k\delta+\delta^{\prime})$-DP where $\delta^{\prime}>0$, and not $(\epsilon^{\prime},\delta)$-DP as claimed.
@@ 3.1.4 Privacy proof:
To address these issues, we propose a new proof of the Differential Privacy guarantees for our algorithm in Theorem 3.2, whose proof is in Appendix A.2. This proof clarifies the ambiguities or errors present in the previous works, and ensures that our privacy guarantees are met.
@@ 3.1.5 Improved convergence bound:
iterations, with probability at least $1-\tau^{-\Omega(p+1-q)}-e^{-\Omega(d)}$, we have
@@
Note: This bound is at least as tight as the one proposed in (Balcan et al., 2016). Indeed the following bound holds:
[…]
which is tight only if $\forall j,\|\bm{X}_{ij}^{l}\|=\max_{j}\|\bm{X}_{:j}^{l}\|$. and can be in practice much tighter depending on $\bm{A}$. Our privacy proof (given in Theorem 3.2) does not restrict us to have $\epsilon\leq 1$. We provide the proof in Appendix A.2.
@@ 4 Analysis of Runtime-independent bounds
The runtime-independent bound that we develop in this section and present in Theorem 4.1 is a combination of two bounds. The first applies for matrices with large coherences, and is independent of the matrix coherence parameter. The second applies to matrices with low $\mu_{1}$-coherence, but the bound is dependent on the $\mu_{1}$-coherence of the matrix. We also show that our method achieves a bound which is at least as tight as both proposed bounds.
@@ 4.1 Proposed runtime-independent convergence bound
The proposed runtime-dependent bound that we presented earlier (Theorem 3.4) involves $\max_{i}\|\bm{X}_{i:}\|_{2}$, which depends on values that change during the execution of the algorithm. Similarly to (Hardt & Price, 2014; Hardt & Roth, 2013), we now propose a runtime-independent convergence bound for our method. This bound, presented in Theorem 4.1, is a combination of two bounds, one that is suited for matrices with high $\mu_{1}(\bm{A})$, and another that applies for matrices with small $\mu_{1}(\bm{A})$. We provide its proof in Appendix A.3. We note that for a matrix $\bm{A}$, the proposed bound is tighter than the one proposed in (Balcan et al., 2016) when $\min(1,\mu_{1}(\bm{A}))\leq\sqrt{p\cdot\log({d})}\cdot\mu_{0}(\bm{A})$. This is likely to happen in practice. We highlight this fact in Section 5.1.
@@
5 Empirical comparison of the proposed bounds
In previous sections, we proposed both runtime-dependent and runtime-independent bounds of convergence for our algorithm. The runtime-dependent bound is dependent on the matrix of interest $\bm{A}$ and on the matrices $\bm{X}_{l}$ that are computed iteratively. The runtime-independent bound (4.1) is dependent only on the matrix of interest $\bm{A}$. We now provide a few examples to justify the usefulness of the proposed bounds.
5.1 Empirical comparison of the runtime-independent bounds
As explained in Section 4.1, the proposed runtime-independent bound is tighter than those proposed in (Hardt & Price, 2014; Balcan et al., 2016) when $\min(1,\mu_{1}(\bm{A}))\leq\mu_{0}(\bm{A})\sqrt{p\cdot\log({d})}$, where $\bm{U}$ denotes the eigenvectors of $\bm{A}$. We now provide an illustration of how the bounds compare for a practically relevant example, namely recommender systems.
Table 1: Statistics of datasets
Dataset Users Items Interactions $\mu_{0}$ $\mu_{1}$ $\Delta_{l}^{a}$ (existing) ${\Delta_{l}^{b}}$ (ours)
Gowalla 29,858 40,981 1,027,370 0.36 82.48 1.08 ×$\times\sqrt{p}$ 1
Yelp2018 31,668 38,048 1,561,406 0.40 60.42 1.29 ×$\times\sqrt{p}$ 1
Amazon-book 52,643 91,599 2,984,108 0.33 52.95 1.10 ×$\times\sqrt{p}$ 1
5.1.1 Recommender systems:
To compute several parameters of SOTA recommender system models, we need to compute leading eigenvectors of a matrix denoted as “the normalized item-item matrix”. This is equivalent as computing the right singular vectors of “the normalized adjacency matrix” (Shen et al., 2021). The most widely used datasets in recommender systems research are Gowalla (Cho et al., 2011), Yelp2018 (He et al., 2020) and Amazon-Book (Wang et al., 2019). Table 1 presents key statistics of these datasets, and compares the sensitivity bounds for the task of computing the leading eigenvectors of interest. As we can deduce from Table 1, the proposed bound allows us to converge to solutions with much smaller $\eta$ values, especially when the desired number of factors $p$ is large.
5.2 Empirical comparison of the runtime-dependent bounds
[…]
$r(k,n)=\frac{\mathbb{E}[\Delta_{0}^{a}]}{\mathbb{E}[\Delta_{0}^{b}]}.$ (18)
5.2.1 Theoretical approximation of $r(k,n)$:
We can find an asymptotic approximation of $r(k,n)$ using Approximation 5.1. The derivation is provided in Appendix A.4.
Approximation 5.1.
[…]
$\mathbb{E}[\Delta_{l}^{a}]^{2}\approx\left(\sqrt{2\sigma^{2}\cdot\log(kn)\frac{k^{2}}{n^{2}}}+\frac{k\mu}{n}\right)$ (19)
$\mathbb{E}[\Delta_{l}^{b}]^{2}\approx\left(\sqrt{2\sigma^{2}\cdot\log(n)\frac{k}{n^{2}}}+\frac{k\mu}{n}\right).$ (20)
5.2.2 Empirical approximation of $r(k,n)$:
We now want to compare theoretically and empirically how tight our proposed bound ($\Delta_{0}^{b}$) is compared to the previous one ($\Delta_{0}^{a}$). The scaling of the noise at the first iteration of the algorithm depends on the random initialization of the eigenvector matrix $\bm{X}_{0}$, which does not depend on the matrix of interest $\bm{A}$. We can therefore empirically estimate $\mathbb{E}[\Delta_{0}^{a}]$ and $\mathbb{E}[\Delta_{0}^{b}]$ on randomly initialized matrices $\bm{X}_{0}$ to compare the tightness of both bounds.
Experiment: $r(k,n)$ depends on the number of factors $k$ and on $n$, where ($n\times n$) is the size of $\bm{A}$. We therefore seek to compute it for multiple values of $n$ ($n\in\{8000,12000\}$) and a range of values for $k$ (between 64 and 4000, with steps of 64), for each value of $n$. We use Theorem 5.1 to approximate it theoretically. To measure it empirically, we compute $\Delta_{0}^{a}$ and $\Delta_{0}^{b}$ based on $\bm{X}_{0}$ in the first step of the algorithm (it is independent of the matrix of interest $\bm{A}$), for $t$ runs of the algorithm. We can then use the $t$ measures to estimate $\mathbb{E}[\Delta_{0}^{a}]$ and $\mathbb{E}[\Delta_{0}^{b}]$. We claim that $t=5$ is enough to see a general tendancy, as we have estimates for many ($k,n$) couples. We can also compare the empirical estimates to the proposed asymptotic approximations specified in Approximation 5.1.
We give our results in Figures 1 and 2. We can see empirically and theoretically that the proposed noise scaling $\Delta_{l}^{b}$ is much tighter than $\Delta_{l}^{a}$ at the first step of the algorithm. Our theoretical approximation is conservative and underestimates how much tighter the bound is initially, compared to what we observe in practice. The proposed bound is tighter by a multiplicative factor on the first step and can therefore drastically reduce the impact of the noise introduced by DP on the convergence of the Privacy Preserving power method, as the algorithm is usually run for very few steps ($L$ is usually in the range of $1$-$5$).

### v2

@@ Abstract
The randomized power method has gained significant interest due to its simplicity and efficient handling of large-scale spectral analysis and recommendation tasks. However, its application to large datasets containing personal information (e.g. web interactions, search history, personal tastes) raises critical privacy concerns. This paper addresses these issues by proposing enhanced privacy-preserving variants of the method, which reduce the variance of the noise introduced to achieve Differential Privacy (DP). We also adapt the method to a decentralized framework with a low computational and communication overhead, while preserving the accuracy. We leverage Secure Aggregation to run the algorithm on data distributed among multiple users or devices, while using a noise scale in the decentralized setting that is similar to the one in the centralized setting yet without revealing individual data. We improve upon existing convergence bounds for both the centralized and decentralized versions.
@@ 1 Introduction
The randomized power method is a simple and computationally efficient tool for solving large-scale linear algebra problems, particularly in spectral analysis and recommendation systems. It approximates the principal singular vectors of a matrix with low computational complexity (nearly linear), making it particularly suitable for analyzing large modern datasets.
@@
To address and quantify privacy leakages, Differential Privacy (DP) has emerged as a powerful mathematical framework that provides strong formal guarantees and mitigate potential privacy leaks of an algorithm. DP ensures that the output of an algorithm reveals little about any individual record in the input. Several works have attempted to apply DP to the randomized power method. For instance, Hardt & Price (2014) and Balcan et al. (2016) introduced differentially private versions of the method, and Guo et al. (2021); Wang & Xu (2020) explored federated approaches to private power iteration. Other close works explore centralized differentially private versions of PCA (Liu et al., 2022) and federated (Wang & Xu, 2020; Briguglio et al., 2023) and give optimal convergence bounds given certain assumptions.
@@ 1.1 Contributions
We propose an improved differentially private version of the randomized power method, with tighter convergence bounds, reducing the noise required to achieve privacy and therefore improving the accuracy of the method. In the process of deriving these bounds, we present a new, more straightforward proof of a privacy result for the differentially private randomized power method. [^1] This proof addresses some minor mistakes in a privacy proof from Hardt & Roth (2012, 2013) that have been reproduced in several follow-up works (Hardt & Price, 2014; Balcan et al., 2016). The modified result allows us to use a wider range of privacy parameters.
@@ 2.2.1 Approximate differential privacy:
A randomized algorithm $\mathcal{M}$ provides ($\epsilon,\delta$)-differential privacy if for all datasets $D_{1}$ and $D_{2}$ differing on at most one element (adjacent datasets), and for all measurable subsets $S\subseteq\text{Range}(\mathcal{M})$, the following holds:
@@ 2.2.2 $\ell_{2}$-sensitivity:
where $\|\cdot\|_{2}$ represents the $\ell_{2}$-norm.
@@ 2.2.3 Zero-Concentrated Differential Privacy (zCDP):
where $D_{\alpha}$ is the Rényi divergence of order $\alpha$ and $\rho$ is a positive parameter controlling the trade-off between privacy and accuracy (smaller values of $\rho$ imply stronger privacy guarantees).
@@ 2.3 Privacy preserving power iteration:
Let $k$ denote the target rank, $q$ a small real and $\eta$ a small positive value controlling the precision of the approximation. The goal of the randomized power method is to compute a matrix $\bm{Q}\in\mathbb{R}^{n\times p}$, where $p=k+q$, such that the range of $\bm{Q}$ approximates the range of the top $k$ eigenvectors of $\bm{A}$, which are denoted $\bm{U}_{k}$. Specifically, we aim to ensure that:
$\|\bm{U}_{k}-\bm{Q}\bm{Q}^{\top}\bm{U}_{k}\|\leq\eta.$
The privacy-preserving randomized power method extends this approach to protect sensitive information in $\bm{A}$ during computation. Prior work (Hardt & Price, 2014; Balcan et al., 2016) show that under approximate Differential Privacy (DP) parameters $(\epsilon,\delta)$, $\bm{Q}$ can be computed using Algorithm 1, where the sensitivity $\Delta_{l}$ at iteration $l$ is bounded by $\Delta_{l}^{\text{prior}}\triangleq\sqrt{p}\cdot\|\bm{X}_{l}\|_{\max}$. Here, $\Delta_{l}^{\text{prior}}$ represents an $\ell_{2}$-sensitivity bound that quantifies the impact of a single-element change in $\bm{A}$ on $\bm{A}\bm{X}^{l-1}$. Specifically, $\Delta_{l}\leq\Delta_{l}^{\text{prior}}$.
With this bound [^2] As defined in Fig. 3 of (Hardt & Price, 2014), Alg. 2 of (Balcan et al., 2016) and Theorem 6 of (Guo et al., 2021)., the standard deviation of the noise $\sigma_{l}$ associated to DP at iteration $l$ is given by:
$=\Delta_{l}^{\text{prior}}\cdot\epsilon^{-1}\sqrt{4L\log(1/\delta)}.$ (7)
As a first contribution, we refine this sensitivity analysis by introducing a significantly tighter bound for $\Delta_{l}$, reducing the noise required for DP and therefore allowing tighter convergence bounds (smaller $\eta$).
@@ 2.4 Existing convergence bounds
To our knowledge, the strongest bound for the randomized power method as defined in Algorithm 1 with $\Delta_{l}=\Delta_{l}^{prior}\triangleq\sqrt{p}\cdot\|\bm{X}_{l}\|_{\max}$ is given in Balcan et al. (2016). In Theorem 2.4, we present a reduction of their Theorem 3.1. The theorem, and several others we present relies on the verification of the conditions in Assumption 2.3.
@@
Let $\bm{A}\in\mathbb{R}^{d\times d}$ be a symmetric data matrix. Fix target rank $k$, intermediate rank $q\geq k$, and iteration rank $p$, with $2q\leq p\leq d$. Suppose the number of iterations $L$ is set as $L=\Theta(\frac{\sigma_{k}}{\sigma_{k}-\sigma_{q+1}}\log(d))$. Let $\epsilon,\delta\in(0,1)$ be the differential privacy parameters. Let $\bm{U}_{q}\in\mathbb{R}^{d\times q}$ be the top-$q$ eigenvectors of $\bm{A}$ and let $\sigma_{1}\geq\cdots\geq\sigma_{d}\geq 0$ denote its eigenvalues. Then if Assumption 2.3 holds, Algorithm 1 with $\Delta_{l}=\Delta_{l}^{prior}$ gives us the following runtime-dependent convergence bound:
@@ 3.1.1 Adjacency notion:
In prior works (Hardt & Price, 2014; Balcan et al., 2016), two matrices $\bm{A}$ and $\bm{A}^{\prime}$ are considered adjacent (denoted $\bm{A}\sim\bm{A}^{\prime}$) if they differ by a change in a single element, with a Frobenius norm of at most 1:
$\bm{A}^{\prime}=\bm{A}+c\cdot\bm{e}_{i}\bm{e}_{j}^{\top},$ (11)
where $\bm{e}_{i},\bm{e}_{j}\in\mathbb{R}^{n}$ are canonical basis vectors of dimension $d$, $c\leq 1$ represents the magnitude of the change, and $0\leq i,j<n$. This adjacency notion models a sensitive change as a modification to a single element in $\bm{A}$ to protect individual element-wise updates under Differential Privacy. However, since $\bm{A}$ is symmetric positive semi-definite and changes must preserve this property, this necessitates updates on the diagonal elements. We propose a new, more general, notion of adjacency to allow for other types of updates:
@@
where $\bm{C}\in\mathbb{R}^{n\times n}$ is a symmetric matrix representing the update, subject to the condition:
$\sqrt{\sum_{i=1}^{n}\|\bm{C}_{i,:}\|_{1}^{2}}\leq 1.$ (13)
The proposed adjacency notion is strictly more general than Equation (11). For example, setting $\bm{C}=c\cdot\bm{e}_{i}\bm{e}_{j}^{\top}$ recovers the original definition. Additionally, $\bm{C}$ can have non-zero entries on the diagonal, anti-diagonal, or any symmetric pattern, allowing for a variety of updates while maintaining symmetry.
In the context of recommender systems where $\bm{A}=\bm{R}^{\top}\bm{R}$ and $\bm{R}$ is a user-item interaction matrix, this adjacency notion allows for element-wise changes in $\bm{R}$ (i.e. protect user-item interactions). By adopting this more general adjacency definition, we make our privacy guarantees applicable to a wider range of real-world scenarios.
@@ 3.1.2 Improved sensitivity bound:
The previous sensitivity bound $\Delta_{l}^{\text{prior}}$ (Hardt & Price, 2014; Balcan et al., 2016) for $\bm{A}\bm{X}^{l-1}$ was estimated by extrapolating from the case where $\bm{X}^{l-1}\in\mathbb{R}^{n\times 1}$ to the general case $\bm{X}^{l-1}\in\mathbb{R}^{n\times p}$. This leads to an overestimation of the sensitivity and to unnecessarily large noise addition. We directly analyze the change in $\bm{X}^{l-1}\in\mathbb{R}^{n\times p}$ and derive a tighter bound on the sensitivity $\Delta_{l}$. By using our proposed adjacency notion from Equation (12), we establish the following result (proof provided in Appendix A.1):
@@
Let $\bm{A}^{\prime}$ be defined as in Equation (12), and consider the sensitivity $\Delta_{l}=\|\bm{A}^{\prime}\bm{X}^{l}-\bm{A}\bm{X}^{l}\|_{F}$. Then, $\Delta_{l}$ is upped-bounded as follows:
$\Delta_{l}\leq\max_{i}\|\bm{X}_{i,:}^{l}\|_{2}\triangleq\hat{\Delta}_{l}.$ (14)
This improved bound $\hat{\Delta}_{l}$ is generally tighter than the previous bound $\Delta_{l}^{\text{prior}}=\sqrt{p}\cdot\|\bm{X}_{l}\|_{\max}$, as it eliminates the dependence on $\sqrt{p}$ and directly relates to the maximum row norms of $\bm{X}^{l}$. By reducing the estimated sensitivity, we can add less noise while still satisfying Differential Privacy, thereby improving the utility of the algorithm.
@@ 3.1.3 Privacy proof:
To address these issues, we propose a result with a new proof of the Differential Privacy guarantees for our algorithm in Theorem 3.2, whose proof is in Appendix A.2.2. This proof clarifies the ambiguities or errors present in the previous works (see Appendix A.2), and ensures that our privacy guarantees are met.
@@ 3.1.4 Improved convergence bound:
iterations, with probability at least $1-\tau^{-\Omega(p+1-q)}-e^{-\Omega(d)}$, we have:
@@
Note: This bound is at least as tight as the one proposed in (Balcan et al., 2016), and can be in practice much tighter depending on $\bm{A}$. Our privacy proof (given in Theorem 3.2) does not restrict us to have $\epsilon\leq 1$. We provide the proof in Appendix A.2.2.
@@ 4.1 Proposed runtime-independent convergence bound
We presented in Theorem 3.4 a runtime-dependent convergence bound involving $\max_{i}\|\bm{X}_{i,:}\|_{2}$, which is only known during the execution of the algorithm. To provide a more practical analysis, we now derive a runtime-independent convergence bound in Theorem 4.1. This bound is a combination of two cases:
  • High Coherence Case: Applies to matrices with large coherence parameters, and the bound is independent of the matrix’s coherence.
  • Low Coherence Case: Applies to matrices with low $\mu_{1}$-coherence, with the bound depending explicitly on $\mu_{1}(\bm{A})$.
We demonstrate that our method achieves a convergence bound that is at least as tight as both of these bounds.
We note that for a matrix $\bm{A}$, the proposed bound is tighter than the one proposed in (Balcan et al., 2016) when $\min(1,\mu_{1}(\bm{A}))\leq\sqrt{p\cdot\log({d})}\cdot\mu_{0}(\bm{A})$. This is likely to happen in practice, as we highlight in Section 6.1.1.